Can you please help me with this fraction problem?
15/16b = 5/12
The "b" is next to the fraction line in between the 15 & 16

Found 2 solutions by AmparoV, richwmiller:
Answer by AmparoV(10)

15/16b=5/12
You would times both sides by 15/16 to get the variable by itself. You would be left with b= 25/64 as your answer.
Hope this helps.

Answer by richwmiller(17219)